OpenStudy (anonymous):

what is 7 times the square root of 96m^3

OpenStudy (jdoe0001):

can you SIMPLIFY 96?

OpenStudy (anonymous):

16 and 6

OpenStudy (jdoe0001):

and would you say \(16=4^2\)? so really \(\sqrt{4^2\times6m^3}\)?

OpenStudy (anonymous):

Nevermind I think I got it as \[28m \sqrt{6m}\]

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

:)